OpenStudy (anonymous):

Determine the numbers of atoms in each of the following 5.40 g B 0.250 mol k 0.0384 mol k 0.02550 g pt 1.00 x 10^-10 g Au

OpenStudy (anonymous):

Boron - (5.40 / 10.811) * 6.022x10^23 = 3.01x10^23 atoms K - 0.25 * 6.022x10^23 = 1.51x10^23 atoms K - 0.0384 * 6.022x10^23 = 2.312x10^22 atoms Platinum - (0.02550 / 195.08) * 6.022x10^23 = 7.8717x10^19 atoms Gold - (1.00x10^-10 / 196.97)* 6.022x10^23 = 3.06 x 10^11 atoms